OpenStudy (anonymous):

using Limit Comparison Test, is the summation of 1/((n^3)(sin n)^2) from n equals 1 to infinity convergent or divergent?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \frac{ 1 }{ n^3 \sin ^2 n}\]

OpenStudy (anonymous):

convergent

OpenStudy (anonymous):

divergent..

OpenStudy (anonymous):

How did you arrive at your answers? Please help me solve.

OpenStudy (anonymous):

use 1/n3 then u get 1/sin2n which isa positive term hence as 1/n3 converges it tooooooooo

OpenStudy (anonymous):

but 1/sin^2n does not converge..

OpenStudy (anonymous):

I mean limit does not exist..

OpenStudy (anonymous):

there is no need that the outcome of the test is converging it should be just positive

OpenStudy (anonymous):

but limit must exist

OpenStudy (anonymous):

we are taking n as natural that is a multiple of 57 degrees so it is positive and it is enough to be +ve for that test

OpenStudy (anonymous):

what is the sum then..

OpenStudy (anonymous):

test doesnt say the sum we use it to just explain convergence of such complex series whose cant be determined

OpenStudy (anonymous):

easily

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Limit_comparison_test you should get limit point such as c, but lim_>inf sin^2n does not exist..

OpenStudy (anonymous):

Guys, accdg to the ratio test, this is inconclusive. What can you say about this? :(

OpenStudy (anonymous):

THERE is no need that a converging series should follow all the tests 'if it satisfies any one it is enough